# NoProp: Training Neural Networks without Back-propagation or Forward-propagation

### v1

@@ 5 Conclusion
Using the denoising score matching approach that underlies diffusion models, we have proposed NoProp, a forward- and back-propagation-free approach for training neural networks. The method enables each layer of the neural network to be trained independently to predict the target label given a noisy label and the training input, while at inference time, each layer takes the noisy label produced by the previous layer and denoises it by taking a step towards the label it predicts. We show experimentally that NoProp is significantly more performant than previous back-propagation-free methods while at the same time being simpler, more robust, and more computationally efficient. We believe that this perspective of training neural networks via denoising score matching opens up new possibilities for training deep learning models without back-propagation, and we hope that our work will inspire further research in this direction.
@@ Appendix B Trainable noise schedule for NoProp-CT
Figure 6: Models used for training when the class embedding dimension is different from the image dimension. Left: model for the discrete-time case. Right: model for the continuous-time case. conv: convolutional layer. FC: fully connected layer (number in parentheses indicates units). concat: concatenation. pos emb: positional embedding (number in parentheses indicates time embedding dimension). When the class embedding dimension matches the image dimension, the noised label and the image are processed in the same way before concatenation in each model. Note that batch normalization is not included in the continuous-time model.

### v2

@@ 5 Conclusion
Using the denoising score matching approach that underlies diffusion models, we have proposed NoProp, a novel approach for training neural networks without full forward- or back-propagation across network blocks. The method enables each block of the neural network to be trained independently to predict the target label given a noisy label and the training input, while at inference time, each block takes the noisy label produced by the previous block and denoises it by taking a step towards the label it predicts. Our experiments show that NoProp is a viable learning algorithm, achieving performance comparable to full back-propagation on the same architecture, while offering advantages in biological plausibility and enabling parallel training. However, our experiments focus on relatively small-scale classification datasets, and evaluating NoProp on larger datasets with higher-resolution inputs will be an important direction for future work. In particular, scaling will require addressing the cost of broadcasting high-dimensional inputs to all blocks and potentially optimizing communication strategies between blocks while preserving locality. While we have applied NoProp only to classification tasks, the underlying framework of block-wise local training may be useful in broader settings. For example, in continual learning, the lack of gradient flow across blocks might help localize task interference, potentially reducing catastrophic forgetting. Although we do not investigate this directly, we highlight it as a possible direction for future research. We believe that the perspective of training neural networks via denoising score matching opens up new possibilities for training deep learning models without full back-propagation, and we hope that our work will inspire further research in this direction.
Acknowledgments
Qinyu Li is supported by the Oxford-Radcliffe Graduate Scholarship and the EPSRC CDT in Modern Statistics and Statistical Machine Learning (EP/S023151/1).
@@ Appendix B Trainable noise schedule for NoProp-CT
Figure 5: Models used for training when the class embedding dimension is different from the image dimension. Left: model for the discrete-time case. For Tiny ImageNet, an additional convolutional layer with 256 output channels, followed by ReLU, 2×2 max pooling, and dropout (p = 0.2), is inserted before concatenation. Right: model for the continuous-time case. conv: convolutional layer. FC: fully connected layer (number in parentheses indicates units). concat: concatenation. pos emb: positional embedding (number in parentheses indicates time embedding dimension). When the class embedding dimension matches the image dimension, the noised label and the image are processed in the same way before concatenation in each model. Note that batch normalization is not included in the continuous-time model.
Dataset One-hot Prototype Label dim Params
Discrete-time
MNIST Yes No 10 0.92M
No No 20 0.92M
No Yes 784 1.40M
CIFAR-10 Yes No 10 1.22M
No No 20 1.22M
No Yes 3072 1.99M
CIFAR-100 Yes No 100 1.25M
No No 20 1.23M
No Yes 3072 2.00M
Tiny ImageNet Yes No 200 2.45M
No Yes 12288 4.30M
Continuous-time
MNIST Yes No 10 0.87M
No No 20 0.87M
No Yes 784 1.48M
CIFAR-10 Yes No 10 1.16M
No No 20 1.17M
No Yes 3072 2.07M
CIFAR-100 Yes No 100 1.20M
No No 20 1.18M
No Yes 3072 2.08M
Table 5: Model parameters.